# BCDNet: A Convolutional Neural Network For Breast Cancer Detection thanks: 2the corresponding author

### v2

@@ Abstract
Previous research has established that breast cancer is a prevalent cancer type, with Invasive Ductal Carcinoma (IDC) being the most common subtype. The incidence of this dangerous cancer continues to rise, making accurate and rapid diagnosis, particularly in the early stages, critically important. While modern Computer-Aided Diagnosis (CAD) systems can address most cases, medical professionals still face challenges in using them in the field without powerful computing resources. In this paper, we propose a novel CNN model called BCDNet, which effectively detects IDC in histopathological images with an accuracy of up to 89.5% and reduces training time effectively. The code is available at https://github.com/YujiaLin-523/BCDNet
@@ I Introduction
Despite the availability of various diagnostic methods such as mammography, digital breast tomosynthesis, breast ultrasound, and magnetic resonance imaging, accurately and rapidly diagnosing breast cancer and IDC remains a significant challenge for radiologists and other medical staff due to the increasing number of patients and the complexity of the diagnostic process. To assist pathologists, Computer-Aided Diagnosis (CAD) systems have been developed as complementary tools to conventional diagnostic methods [4, 5]. These systems analyze medical images or other data and provide diagnostic results to pathologists, specifically aiding in the diagnosis of IDC by saving time while maintaining accuracy. The continuous advancements in Deep Learning, especially in Computer Vision, contributes a lot to the improvement of CAD. In Medical Image Processing, CNN-based methods are widely used for breast cancer detection due to CNN’ s ability to learn features automatically rather than relying on hand-crafted features. However, current CNN models are often challenging to train due to their relatively large number of parameters and the limited computing resources in some medical institutions. This limitation hampers the ability to quickly adapt to changes in datasets, such as the emergence of a new type of cancer. To address this, we developed BCDNet that is both efficient and easy to train, enabling deployment on edge devices and rapid modification for new datasets. Our main contributions can be concluded as follows:
  • We propose a novel architecture BCDNet, which is efficient and easy to train.
  • We prove the superiority of BCDNet by comparing it with ResNet 50 and ViT-B-16.
@@ III Method
Figure 1: The architecture of BCDNet. We combine Convolutional layers, Pooling layers and ReLU with BatchNorm to build this model.
The architecture of BCDNet is shown in Fig. 1. The model consists of a series of convolutional layers, pooling layers, activation layers, fully-connected layers, Batch Normalization layers and Dropout layers. This structured combination of layers is designed to efficiently and accurately detect IDC in histopathological images. In the following sections, we will delve into the specific configurations and functionalities of each layer type within BCDNet.
III-A Convolutional Layer
Convolutional Neural Networks (CNNs) employ a mathematical operation called convolution for the purpose of feature extraction [10, 11, 12]. This operation, a fundamental aspect of CNNs, is simply depicted in the following equation:
$F=I*K$ (1)
where $F$ represents the feature map, $I$ is the input, and $K$ is the kernel. The kernel, a matrix of weights, is paramount in extracting significant features from the input, thereby enabling the network to learn and identify various patterns. The kernel’s weights are optimized using backpropagation to minimize the loss function. This convolution process can also be illustrated in detailed by
$s(t)=(x*w)(t)=\sum_{a=-\infty}^{+\infty}x(a)w(t-a)$ (2)
Feature maps in CNN are generated by convolving a kernel with the input data. CNNs employ parameter sharing, where the same kernel weights are reused across different input regions, significantly reducing the total number of parameters needed. This approach efficiently captures complex patterns and features across various images while maintaining a simpler model structure.
III-B Pooling Layer
Pooling is an essential down-sampling method used in CNN to enhance invariance and combine semantically similar features, thus ensuring robust feature extraction [10]. The primary goal of pooling is to reduce the spatial dimensions of a feature map, thereby decreasing the amount of data while preserving crucial features for the learning process. Max pooling, a common pooling operation, selects the largest value within a designated window or kernel area, effectively reducing the dimensionality of the feature map and making the representation less sensitive to minor shifts or alterations in the input.
In mathematical terms, consider a feature map $F$ of dimensions $i\times j$, and a pooling window (or kernel) $P$ of size $k\times k$. Max pooling can be expressed as:
$M(i,j)=\max_{0<p,q\leq k}F_{i+p,j+q}$ (3)
where $M$ is the down-sampled feature map, and $M_{(i,j)}$ is the maximum value within the $k\times k$ window centered at $(i,j)$ in the original feature map. Max pooling, a fundamental technique in convolutional neural networks (CNNs), involves sliding a pooling window over the feature map and selecting the maximum value within each region. This process effectively reduces the dimensionality of the feature map while retaining its most salient features, thereby enhancing the network’s efficiency and ability to extract invariant features. In summary, max pooling is crucial in CNNs for down-sampling, enhancing invariance, and pinpointing the most prominent features within the input data.
III-C Activation Layer
The Rectified Linear Unit (ReLU), is a widely utilized activation function in neural networks [13]. The function is defined as the following equation and can be seen in, equates values less than zero to zero, and values greater than or equal to zero to passed in value.
$f(x)=max(0,x)=\begin{cases}0,x\leq 0\\ x,x>0\end{cases}$ (4)
ReLU introduces non-linearity into neural networks without compromising computational efficiency, avoiding the vanishing gradient problem seen in sigmoid and tanh functions. This property allows neural networks to learn complex patterns more effectively. Additionally, ReLU prevents gradient saturation, facilitating faster and more efficient training. ReLU is critical in modern neural networks for its role in enhancing non-linearity, improving computational speed, and aiding in model convergence.
III-D Fully-Connected Layer
The Fully-Connected (FC) layer can be viewed as performing a mapping of all the features in a given feature space [10, 11, 12]. Its task is to transform these features through a combination of linear operations followed by non-linear activation functions, enabling complex decision-making processes. In CNN, the FC layer processes the features extracted by the convolutional and pooling layers and maps them to the final decision output. This can be expressed mathematically as:
$y=f(Wx+b)$ (5)
where x represents the input features, $W$ is the weight matrix, $b$ is the bias term, $f$ is the activation function ReLU in our work, and $y$ is the output. In the upper representation, $Wx$ denotes a linear transformation that maps the input vector $x$ into a new space. The weight matrix $W$ in an FC layer assigns a unique set of weights to each row, combining input features to compute each output neuron’s value. The bias b adjusts this mapping, increasing the model’s flexibility. This mapping transforms processed features into new representations, suitable for final decision-making tasks. This integration and transformation of features in the FC layer are crucial for supporting classification, regression, and other decision-making tasks in neural networks.
III-E Batch Normalization Layer
Batch Normalization enhances model performance and stability by normalizing layer inputs during training, mitigating issues like exploding gradients. Additionally, it can accelerate convergence since it transforms inputs into a certain distribution. In CNN, Batch Normalization plays a crucial role by normalizing the outputs of convolutional layers, ensuring that each layer receives inputs with a consistent scale. This process helps stabilize and speed up the training of deep CNNs. By reducing internal covariate shift, Batch Normalization allows for higher learning rates, thereby improving training efficiency and model performance [14]. Given input x, this process is:
@@ III-E Batch Normalization Layer
III-F Dropout
Dropout is a regularization technique used to prevent overfitting in neural networks [15]. By randomly setting a fraction p of the neurons’ activations to zero during the training, the network is encouraged to learn more robust features, thereby improving its generalization capabilities. The dropout process can be expressed as:
@@ IV Experiments
To evaluate BCDNet’s performance, we conducted a comprehensive experiment between BCDNet, ResNet50 and ViT-B-16, which are renowned for their robust performance in image classification tasks. ResNet50, a deep residual network, is celebrated for its pioneering use of residual learning to solve vanishing gradient problems [16], while ViT-B-16, a Vision Transformer, represents a shift towards transformer architectures in computer vision, demonstrating strong performance by leveraging self-attention mechanisms [17]. Training was conducted on a server with 2 NVIDIA RTX 4090 GPUs, leveraging a cross-entropy loss function and the Adam optimizer. The learning rate was initially set to 0.005, with a StepLR scheduler to facilitate convergence. Our study is based on two open-source IDC image datasets, and we assess models on accuracy, training time, and memory consumption. The results are summarized in Fig 2 and Table I.
Figure 2: The accuracy and training time of BCDNet, ResNet 50 and ViT-B-16 on IDC regular dataset.
Figure 3: The accuracy and training time of BCDNet, ResNet 50 and ViT-B-16 on BreaKHis v1 dataset.
@@ IV-A Dataset
In this study, we firstly utilized an Invasive Ductal Carcinoma (IDC) image dataset from kaggle, which is termed as IDC regular by us in this paper. IDC regular dataset comprises histopathological scans of breast cancer tissues, annotated to indicate the presence or absence of IDC. Meanwhile, we also use BreaKHis v1 dataset, which is composed of multiple types of breast cancer histopathology images obtained under various imaging conditions. The two datasets are divided into training, validation and testing sets by a proportion of 7:2:1 respectively to enable the evaluation of our proposed model, BCDNet, against established benchmarks.
@@ IV-B Data Augmentation
Data augmentation is a strategy designed to expand the size of a dataset, enhancing both the model’s performance and robustness to overfitting [11]. It involves generating new training samples by applying various transformations to the original data.
@@ IV-D Results
Based on Fig 2, 3, 4, 5 and Table I, we can summarize that though BCDNet has a slightly lower accuracy than ResNet 50, it is more efficient in terms of training time and memory consumption even though we used pretrained models of ResNet and ViT in our experiments. BCDNet is more suitable for deployment on edge devices and can be quickly modified for new datasets. Moreover, the convergence speed of BCDNet is also higher. Its loss and accuracy both become stable quite early.
@@
Figure 4: The accuracy and loss curves of the three models on IDC regular dataset. From up to down, the three models are BCDNet, ResNet 50, and ViT-B-16.
Figure 5: The accuracy and loss curves of the three models on BreaKHis v1 dataset. From up to down, the three models are BCDNet, ResNet 50, and ViT-B-16.
@@ V Conclusion
In this paper, BCDNet has shown its advantages that it consumes much less training time and memory consumption while loses a little accuracy, which means it is suitable for the medical system in the impoverished areas or in the wild.
Although we have leveraged a novel architecture to enhance the speed of the model, the accuracy of BCDNet is still not high enough. In the future, the model can be continously improved by using Layer Freezing, which is widely used to minimize the number of parameters. Additionally, Post-Training Quantization (PTQ) can be probably applied to accelerate the whole process when increasing the depth of the model.

### v3

# BCDNet: A Fast Residual Neural Network For Invasive Ductal Carcinoma Detection thanks: 3the corresponding author

@@ Abstract
It is of great significance to diagnose Invasive Ductal Carcinoma (IDC) in early stage, which is the most common subtype of breast cancer. Although the powerful models in the Computer-Aided Diagnosis (CAD) systems provide promising results, it is still difficult to integrate them into other medical devices or use them without sufficient computation resource. In this paper, we propose BCDNet, which firstly upsamples the input image by the residual block and use smaller convolutional block and a special MLP to learn features. BCDNet is proofed to effectively detect IDC in histopathological RGB images with an average accuracy of 91.6% and reduce training consumption effectively compared to ResNet 50 and ViT-B-16.
@@ I Introduction
Despite the availability of various diagnostic methods such as mammography, digital breast tomosynthesis, breast ultrasound, and magnetic resonance imaging, accurately and rapidly diagnosing the large scale of IDC cases remains a significant challenge for radiologists and other medical staff. To assist pathologists, Computer-Aided Diagnosis (CAD) systems have been developed as complementary tools to conventional diagnostic methods [4, 5]. These systems analyze medical images or other data and provide diagnostic results to pathologists, specifically aiding in the diagnosis of IDC by saving time while maintaining accuracy.
Deep Learning has greatly advanced Computer-Aided Diagnosis (CAD) by now. In Medical Image Processing, CNN-based methods are widely adopted for breast cancer detection due to CNNs’ ability to automatically learn features, eliminating the need for hand-crafted features. In detecting IDC, histopathological image features are influenced by factors such as geographical region, ethnicity, and patient age [2], [3]. However, public datasets often fail to fully capture the variety in these images, leading to inaccurate predictions when generalized models are used. Therefore, fine-tuning models on real-world datasets is crucial for improving detection accuracy. At the same time, the demand of integrated and automatic diagnosis system is increasing, requiring the deep learning models to be combined with other software and hardware components to provide a seamless and efficient solution.
Therefore, many existing approaches that use large models is not suitable for this task. A powerful and expensive computer is required to train and deploy such models, which is not always feasible in real-world scenarios. This limitation prevents their practical application in scenarios where fast diagnosis is needed or computing resources are constrained. To address these challenges, we developed BCDNet—a model designed to be both computationally efficient and easy to train. We achieve this by designing the architecture again. In our BCDNet, the input image is firstly upsampled to make it easier to extract features. Then the features are learned by a small convolutional block and a improved MLP. This allows for reducing the parameters and the deployment on edge devices with limited computation ability, enabling fast adaptation to new datasets and making it suitable for practical and real-time medical use. Our main contributions can be concluded as follows:
  • We propose a novel architecture BCDNet, which is efficient and suitable for small-scale datasets.
  • We reassemble the widely used neural network architectures for Breast Cancer Detection to make BCDNet mission-critical.
  • Our model still remains simple and lightweight with almost the same performance as other complex models.
@@ II-A Conventional Diagnosis Methods
II-B Deep Learning-Based Methods
In 2016, Wang et al. [10] introduced a segmentation method to localize potential cancer areas and used GoogLeNet as a backbone for classification. Subsequently, Chougrad et al. [11] proposed a method that enabled several state-of-the-art (SOTA) architectures to outperform traditional ones on the MIAS database, utilizing fine-tuning techniques to adapt ResNet50, VGG16, and Inception v3 for breast cancer detection.
While these methods show excellent performance in laboratory settings, many medical institutions, especially in developing countries, face a lack of computing resources and rich datasets, impeding the clinical application of SOTA models. To address these limitations, we propose BCDNet that is easy to train without rich computing resource.
@@ III Method
Figure 1: The architecture of BCDNet. We combine residual blocks, convolutional blocks, and MLPs to extract features from the input image. The detailed structure of each block is alsoshown in the figure. The blocks are generally composed of convolutional layers, pooling layers, activation layers, normalization layers, and dropout layers in different orders.
To improve the real-time performance and decrease the memory usage while remaining a high accuracy, we propose the architecture of BCDNet, which is composed of a residual block, a convolutional block and a special designed Multi Layer Perceptron (MLP). The residual block will initially upsample the input image so that we can decrease the depth and width of convolutional block to decrease the number of parameters. After the residual block, we use a smaller convolutional block and MLP with Dropout layer to further improve the speed while maintaining the accuracy. In the following sections, we will delve into the specific configurations and functionalities of each layer type within BCDNet.
III-A Residual Block
The residual block is designed for increase the number of the channels of images to learn more necessary and complex features. Since the IDC histopathological images are basically the same, it is of great significance for the model to recognize the tiny differences between them. One common way is to increase the channel of the input images by applying the convolutional layers which will not change the size of the input image. The convolutional layer is the fundamental componet of the modern CNNs [12, 13, 14] and can be represented as:
$Y(i,j)=\sum_{m=1}^{k_{w}}\sum_{n=1}^{k_{h}}X(i+m-1,j+n-1)\cdot K(m,n)$ (1)
where $X(i,j)$ and $Y(i,j)$ are the input and output matrix respectively; $K(m,n)$ is the kernel; $k_{w}$ and $k_{h}$ are the width and height of the kernel, which are equal in most situations.
In the residual block, Batch Normalization is used after the convolutional layers, enabling the model to learn more robust features and boost the speed of convergence [15]. Additionally, it can partly issues like exploding gradients. We simply denote Batch Normalization layer as:
@@ III-A Residual Block
To avoid vanishing gradient problem seen in sigmoid and tanh functions, we choose the Rectified Linear Unit (ReLU) as the activation function in the residual block [16]. The function is defined as the following equation and can be seen in, equates values less than zero to zero, and values greater than or equal to zero to passed in value.
$f(x)=max(0,x)=\begin{cases}0,x\leq 0\\ x,x>0\end{cases}$ (3)
To further improve the performance of BCDNet, we add a residual connection. In 2016, He et al. [17], proposed the Deep Residual Network that contains the residual connection, which is then become the state-of-the-art architecture for image classification. The connection reformulates the layers as learning residual functions with reference to the layer inputs, rather than directly learning unreferenced functions. Applying this shortcut in the neural network allows the model to learn more complex features and decrease the probability of degradation problems as network depth increases.
Mathematically, if $H(x)$ is the desired mapping, the residual block instead learns:
$H(x)=F(x)+x$ (4)
This simple addition allows the network to pass information directly from earlier layers to later layers, thereby facilitating gradient flow during backpropagation and improving the ability to train very deep networks. The residual connections also enable better preservation of low-level features throughout the network, which is beneficial for accurate representation learning. Meanwhile, the original input is upsampled to match the dimensions of the output, which is achived by a $1\times 1$ convolution.
III-B Convolutional Block
By leveraging the residual block, the features in the input image are learned in different channels of the feature maps. However, the features are not extracted and the feature maps are still too large for practical using. Therefore, we build a convolutional block to extract the features from the input image and reduce the spatial dimensions of the feature maps.
The convolutional block is composed of several convolutional layers, pooling layers, activation layers, and normalization layers. The pooling layers here are Max Pooling layers, which can be denoted as:
$M(i,j)=\max_{0<p,q\leq k}F_{i+p,j+q}$ (5)
where $M$ is the down-sampled feature map, and $M_{(i,j)}$ is the maximum value within the $k\times k$ window centered at $(i,j)$ in the original feature map. The Max Pooling layer will preserve the maximum value within the window, which means it can extract the most outstanding features in the feature map. Moreover, the pooling layers can also reduce the spatial dimensions of the feature maps, which is beneficial for reducing the computational complexity and improving the model’s efficiency.
In the convolutional block, we also add several Dropout layers to prevent overfitting and make the model more robust and lightweight [18]. By randomly setting a fraction p of the neurons’ activations to zero during the training, the network is encouraged to learn more robust features, thereby improving its generalization capabilities. The dropout process can be expressed as:
@@ III-B Convolutional Block
III-C Multi-Layer Perceptron (MLP)
Though we regard the Breast Cancer Detection as a binary classification task, we hope that other scholars and medical professionals can also use our model to perform more accurate classification about the certain subtype of Breast Cancer. Thus, we design a more complex MLP than this binary classification task needs. Based on the traditional MLP, we add dropout layers to prevent overfitting and make the model more robust and lightweight. ReLU is also used as the activation function in the MLP, which can make this MLP learn more non-linear relationships between the input features and the output. The structure of this MLP can be seen in Fig 1.
@@ IV Experiments
To evaluate BCDNet’s performance, we conducted a comprehensive experiment between BCDNet, ResNet50 and ViT-B-16, which are renowned for their robust performance in image classification tasks. ResNet50, a deep residual network, is celebrated for its pioneering use of residual learning to solve vanishing gradient problems [17], while ViT-B-16, a Vision Transformer, represents a shift towards transformer architectures in computer vision, demonstrating strong performance by leveraging self-attention mechanisms [19].
Training was conducted on a server with 2 NVIDIA RTX 4090 GPUs, leveraging a cross-entropy loss function and the Adam optimizer. The learning rate was initially set to 0.005, with a StepLR scheduler to facilitate convergence. Our study is based on two open-source IDC image datasets, and we assess models on accuracy, training time, and memory consumption.
Figure 2: The accuracy and training time of BCDNet, ResNet 50 and ViT-B-16. Left: The results of the three models on BreaKHis v1 dataset. Right: The results of the three models on IDC regular dataset.
Figure 3: The curves of the accuracy and loss of BCDNet, ResNet 50 and ViT-B-16 (from the first row to the third row). Left: The curves of the three models on IDC regular dataset. Right: The curves of the three models on BreaKHis v1 dataset.
@@ IV-A Dataset
In this study, we firstly utilized an IDC image dataset from kaggle, which is termed as IDC regular by us in this paper. IDC regular dataset comprises histopathological scans of breast cancer tissues in RGB format, annotated to indicate the presence or absence of IDC. Meanwhile, we also use BreaKHis v1 dataset, which is composed of multiple types of breast cancer histopathology images obtained under various imaging conditions, including different stains, magnifications, and imaging protocols. The two datasets are divided into training, validation and testing sets by a proportion of 7:2:1 respectively to enable the evaluation of our proposed model, BCDNet, against established benchmarks.
@@ IV-D Results
Based on Fig 2, Fig 3 and Table I, we can summarize that though BCDNet has a slightly lower accuracy than ResNet 50, it is more efficient in terms of training time and memory consumption even though we used pretrained models of ResNet and ViT in our experiments. BCDNet is more suitable for deployment on edge devices and can be quickly modified for new datasets. Moreover, the convergence speed of BCDNet is also higher, since its loss and accuracy both become stable quite early.
@@ V Conclusion
In this paper, BCDNet has demonstrated several key advantages, including reduced training time and lower memory consumption, all while maintaining the level of accuracy. Basically, our method can save 12.5% GPU memory than ResNet 50 and 28.8% than ViT-B-16 under our training settings. In terms of time, it is averagely 1.40 times faster than the two baselines on BreaKHis v1 dataset, and 1.35 times faster on IDC regular dataset. The accuracy of BCDNet on two datasets are 93.8% and 89.3% respectively. From our experiments, we can conclude that our pipeline is efficient and effective for IDC detection, which means adding the number of channels and extract features using a shallower convolutional block can accelerate the training process and not decrease the accuracy heavily.